# What Do You Need for Diverse Trajectory Composition in Diffusion Planning?

### v2

@@ A.2.3 Algorithm Hyperparameters
Instead of jointly generating states and actions as in [34], we use a separate inverse dynamics model (as in [1]). This removes the need to tune action versus state weighing in the diffusion objective, and generally performs better than joint prediction [46]. We execute this inverse dynamics model with feedback, executing the action at each time step by giving the model the current state in the plan and the current state from the environment , such that .
A.2.4 Experiment Computing Resources
All experiments for this work were done on a laptop with a 13th Gen Intel Core i9-13950HX and NVIDIA GeForce RTX 4080 Laptop GPU. The final models used to produce the results shown in the paper were trained over roughly 2 weeks, and model inference to generate the final results took roughly 1 week.
A.3 More Environment Details
A.3.1 Grid Environment Details
Our main environment is a small grid where a point agent navigates from one part of the grid to another while avoiding a series obstacles. The observation is the x-y coordinates of the agent, and the action is a x-y step which is normalized to a small step size. In the middle of the maze are several obstacles, which will kill the agent if it stays inside of the block for several steps in a row. Figure 1 shows examples of our environment, including points navigating to and from different locations in the grid. Navigating from one corner of the grid to the other requires steps out of a total allowed episode length of , challenging the diffusion planner’s long-horizon planning. The obstacles also ensure the trajectory must remain compositionally legitimate over the whole horizon, as accumulating error can cause the agent to wander off of the grid to its death.
A.3.2 Grid Environment Possible Trajectories
@@
A.3.3 Data Collection
In the unconditional setting, each trajectory starts in the top-left corner and navigates to the bottom-right corner, randomly switching from moving either downwards or right at each intersection. In total, our unconditional dataset includes unique trajectories out of a possible trajectories.999See Appendix A.3 for proof. This data quantity means our diffusion planner must generalize from a small number of samples to be able to generate samples in the (much larger) space of unseen but legitimate trajectories.
In the conditional setting, we replicate the above data collection process, but for all four corners, to ensure a diverse set of subsequences. We additionally add data starting from each edge intersection point going towards other randomly selected edge intersection points. This ensures that the dataset contains a wide range of subsequences, but that the majority of start-goal states are not seen during training. This dataset contains trajectories.
A.4 Further Architecture Details and Experiments
A.4.1 U-Net is Non-Local and Not Shift Equivariant
U-Net is Not Local: Here we support the claim made in the main text that the U-Net architecture proposed by [34] is not shift equivariant and displays heavy non-locality during the denoising process.
To show this, we follow the procedure of prior works in analyzing diffusion model local field receptiveness by calculating the gradient of an intermediate sample point with respect to the input sequence. We took our final trained U-Net and Eq-Net models in the unconditional environment and generated samples. For each of the denoising steps, we calculated the gradient of the middle pixel in the sequence with respect to the whole input sequence (possible because of auto-differentiation), and averaged these gradients across the generations for each time step. Formally, for a diffusion model where is the next de-noised sample, is the previous sample, and is the denoising step, we calculate for each diffusion time-step from . Because we generate a 1-D sequence, we can plot these gradient maps as an image, where the x-axis is the denoising step and the y-axis represents the sequence gradient.
Figure 9 shows these gradient maps. Visually, it is clear that the U-Net attends early in the denoising process to parts of the sequence far away from the centre pixel, even though the influence is much less strong than the pixels directly surrounding the centre. The size of the receptive field early in training spans far, containing non-zero information as far as the boundary states of the trajectory. On the other hand, Eq-Net has a hard boundary roughly pixels to each side of the centre pixel past which the model cannot attend to in a single denoising step. Like in prior works [36, 52], our diffusion models’ receptive fields progressively shrink during training, starting large and ending small.
We also plot a rough estimate of the local receptive field size at each denoising step. We do this the same way as [52]. We set a threshold value of the th percentile of the value of the absolute gradients in the image, find the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plot the size of that window. Figure 10 shows the results of this, smoothed with a Savitzky-Golay filter. This plot shows that as expected, the local receptive field of the U-Net has a larger maximum value, approaching pixels wide at its max, while the Eq-Net is architecturally capped at .
This is a somewhat surprising result, given that [34] specifically highlights the local receptiveness from convolutions as supporting composition. Given that our results in both the conditional and unconditional setting found that the U-Net architecture struggles with composition, we directly analyzed the network during the diffusion process to gain some insight into whether the model is sufficiently local. We suspect part of the reason for this is because while stride 1 convolutions are locally receptive if applied to the whole sequence, pooling operations and long-stride convolutions both down-sample the image substantially, such that the internal trajectory length at the bottle-neck of the U-Net may be small enough that even a small convolutional kernel is capable of attending to information far across the sequence.
U-Net is Not Shift Equivariant
@@ A.4.1 U-Net is Non-Local and Not Shift Equivariant
A.4.2 Engineering Details for Eq-Net
The main purpose of Eq-Net is not to propose a competitive architecture, but to show a proof of concept that locality and shift equivariance can be baked into architecture to enable stitching. Given this, we did not aggressively tune the architecture parameters of Eq-Net, instead finding values for network depth, kernel size, and channel size that worked well in our benchmark. However, we found it has some limitations and advantages that we will discuss here.
Eq-Net is substantially more memory efficient than the base U-Net model, with the base U-Net having million parameters and Eq-Net only having million parameters.
Despite this, Eq-Net was around half as efficient at both training and inference time. We suspect the former is because the channel sizes grew to be much larger in the intermediate representation layers of the U-Net ( at maximum) while Eq-Net has a fixed channel size of throughout. The latter is likely because Eq-Net is layers deep, which prevents efficient forwards passes because fewer operations can be parallelized.
@@ A.4.2 Engineering Details for Eq-Net
A.4.3 Kernel Size Tuning
@@
Broadly, we found that tuning KER trades off between overall plan consistency (how likely the trajectories were to be locally consistent and stay outside the obstacles) and compositionality. To show this, we trained Eq-Net models in our unconditional setting with 3 KER values: 2, 5, and 12. Our results are shown in Table 3. We found that the intermediate value of 5 generated both a strong diversity of plans as well as reliably well-composed plans, while making the model too non-local made the model too local failed to consistently produce legitimate composed trajectories.
| Large (KER = 2) | Medium (KER = 5) | Small (KER = 12) | |
| Unique Trajectories | |||
| Completion Rate |
We also generated denoising gradient graphs of the small and large kernel Eq-Nets, like in our analysis of U-Net and our final Eq-Net architecture. Figure 13 shows the gradient graphs. As expected, Eq-Net with large kernels is attentive to information across the trajectory, and interestingly produces a gradient map extremely similar to the default U-Net, while Eq-Net with small kernels has an extremely limited receptive field.
Our experiments with the default U-Net architecture and our Eq-Net architectures in total confirm our earlier hypothesis that while CNNs are meant to have strong inductive biases to encourage composition, the particular set of decisions made by [34] means at small data amounts they can still overfit heavily and attend to non-local informatin,
A.4.4 Kernel Diversity
One thing that we found was important when designing Eq-Net was to include both very small (such as k=3) along with the larger kernel layers. When we used only medium sized kernels, it would lead to trajectories that were oddly "ungrounded" to the navigation plane, and would look correct in shape but would be shifted such that the resulting trajectory would nearly always hit obstacles. Figure 14 shows the visual results of this. We found that increasing the kernel size from small to large during training helped, somewhat mirroring the process in normal CNNs where fixed-size kernels stride over the whole image initially (capturing fine detail) and then smaller subsets after each downstride (capturing coarse detail). However, we did not explore whether the exact ordering of the kernel layers in Eq-Net matters or not.
A.4.5 Positional Encoding
When testing how our small receptive field model performs when it has access to positional information for each state, we use the same positional encodings as in [66], with slightly varied parameters. Figure 15 shows a visualization of how we add positional encoding to the Eq-Net architecture. These encodings turn the position of each state into waves forming a geometric progression from which the position of a particular state can be inferred. The encodings are added after the first residual block. Specifically, the positional encodings are defined as
We used a maximum length of instead of used by [66] because of the shorter length of our sequences.
@@ A.4.5 Positional Encoding
A.5 More Experiments
A.5.1 Overfitting Analysis
@@
Here we analyze the effect of overfitting. A common strategy in all machine learning to increase generalization is to perform early stopping, where training is ended before loss on the training set plateaus. Here, we run all of the models shown in the unconditional experiments at gradient increments of and run each model for samples, tracking both their completion rates and number of unique trajectories. Figure 16 shows that overfitting indeed reduces compositionality in the small-data setting, as models early in training achieve both high completion rates (indicating the generated plans are kinematically feasible) as well as more novel trajectories than later in training. However, the models that show stronger compositionality (such as Eq-Net) retain this property late into training (i.e, the losses in compositionality from overfitting plateaus). This hopefully indicates that the proper inductive biases from architecture choices can avoid the overfitting problem.
A.5.2 Does Data Scaling Increase Locality?
In the Results section, we found that scaling data increased the compositionality of our base U-Net model, even though it does not have a strong inductive bias towards locality. To see if the compositionality induced by data scaling is attributable to data increasing locality, we analyzed the models as done previously in Appendix A.4.1 by recording the gradient of the middle pixel for samples across each denoising step for each model. We then plot these directly, and also plot the estimated size of the local receptive field for each model. Following [52], to plot the receptive field we set a threshold value of the th percentile of the value of the absolute gradients in the image, found the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plotted the size of that window.
Figure 17 shows the gradient maps. Note that we name each data scale, with the amount of data corresponding to each name appearing in parentheses (SMM = Small Medium Medium, SM = Small Medium, SSM = Small Small Medium). Visually, it can be seen that the receptive field of the largest model (Full) is smaller than the field of the smallest model (Small) early in the denoising process. The small model attends to information clearly at the bounds of the image around step 20 of the denoising process, while the full data model does not have substantial gradient information past around 150 pixels away from the centre on either end.
Figure 18 shows the plotted local receptive fields, smoothed with a Savitzky-Golay filter. The model trained on the highest data scale has a clearly smaller receptive field compared to the models trained on less data. This indicates that the reason why data scaling may help composition is because it trains the model to have a smaller receptive field, highlighting locality’s importance for composition and the usefulness of our theory. More work should be done to understand why data scaling specifically induces such a property.
A.6 More Discussion
A.6.1 Broader Impacts

### v3

@@ VI-C2 Algorithm Hyperparameters
For Diffusion Planner and CompDiffuser, instead of jointly generating states and actions as in Janner et al. [40], we use a separate inverse dynamics model (as in Ajay et al. [1]). This removes the need to tune action versus state weighing in the diffusion objective, and generally performs better than joint prediction [54]. We execute this inverse dynamics model with feedback, executing the action at each time step by giving the model the current state in the plan and the current state from the environment , such that .
Full details for our Diffusion Planners are given in Table II, Diffusion Policy in Table III, and CompDiffuser in Table IV.
[…]
| Hyperparameter | Navigation | Manipulation |
| Optimizer | AdamW [53] | AdamW [53] |
| Base LR | 2e-4 | 2e-4 |
| Weight Decay | 1e-5 | 1e-5 |
| LR Schedule | Cosine | Cosine |
| Diffusion Prediction | Noise | Clean |
| Sampling Temperature | 0.5 | 0.5 |
| Batch Size | 256 | 64 |
| Convolution Padding | Repeat | Repeat |
| Solver | DDPM | DDPM |
| SDE Form | Continuous | Continuous |
| Diffusion Steps Train | 100 | 100 |
| Diffusion Steps Eval | 500 | 100 |
[…]
| Hyperparameter | Navigation | Manipulation |
| Embedding Dimension | 64 | 64 |
| Layer Sizes | (1,2,2,2) | (1,2,2,2) |
| Planning Horizon | 128 | 512 |
| Model Size | 120 MB | 120 MB |
[…]
| Hyperparameter | Navigation | Manipulation |
| Embedding Dimension | 64 | 64 |
| Model Dimension | 256 | 256 |
| N Heads | 64 | 64 |
| Depth | 8 | 8 |
| Planning Horizon | 128 | 128 |
| Model Size | 74.2 MB | 74.2 MB |
[…]
| Hyperparameter | Navigation | Manipulation |
| Kernel Expansion Rate | 10 | 5 |
| Channel Dimension | 256 | 128 |
| Depth | 15 | 25 |
| Planning Horizon | 128 | 512 |
| Model Size | 61.1 MB | 48.9 MB |
[…]
| Hyperparameter | Maze | Well-Plate | Block-Stack |
| To | 1 | 2 | 6 |
| Tp | 16 | 16 | 16 |
| Ta | 15 | 16 | 16 |
| Img Res | N/A | (160,160) | (160,160) |
| Video Encoder | N/A | ResNet18 | ResNet18 |
| Gradient Steps | 1000000 | 1000000 | 1000000 |
| Pos/Vel Control | Positional | Positional | Positional |
| Architecture | Chi U-Net | Chi U-Net | Chi U-Net |
| Learning Rate | 2e-4 | 1e-4 | 1e-4 |
| Diffusion Prediction | Noise | Noise | Noise |
| Diffusion Steps Train | 100 | 100 | 100 |
| Diffusion Steps Eval | 100 | 100 | 100 |
| Batch Size | 256 | 256 | 256 |
[…]
| Hyperparameter | Navigation | Manipulation |
| High-Level Traj. Horizion | 128 | 512 |
| Single Model Horizon | 32 | 128 |
| # of Trajectories | 5 | 5 |
| Gradient Steps | 1250000 | 1250000 |
| Diffusion Prediction | Clean | Clean |
| Architecture | U-Net | U-Net |
| Batch Size | 250 | 250 |
| Diffusion Steps Train | 100 | 100 |
| Diffusion Steps Eval | 100 | 100 |
| Inference Mode | Parallel | Parallel |
VI-C3 Additional Architecture Experiment Details
Here we show the completion rate of each architecture in our Maze environment. We found that completion dips slightly for Eq-Net near the boundry where it achieves compositionality, and a smaller dip for DiT where it achieves compositionality. See Figure 13. We consistently found that models experienced a dip in completion around the threshold they began to generalize (for Eq-Net around training samples, for DiT around training samples and for U-Net around training samples. Similar behaviour was seen in Pham et al. [67] where diffusion models undergo a three-phase transition as training data scales, starting with memorization, then undergoing a spurious generalization phase where sample quality drops, and finally achieving true generalization where quality is restored along with the achievement of creativity.
VI-C4 Experiment Computing Resources
Experiments used roughly 168 GPU-days of compute cluster use on GPUs equivalent to NVIDIA L40s.
VI-D More Environment Details
Here we describe our environments in more detail.
Lights: We adapt this environment from the Lights Out/Puzzle environment from OgBench [64] with modifications. Specifically, in our environment pressing a button only changes the state of the button itself and not the surrounding buttons, and all trajectories start from the same state where no lights are pressed. We did both of these to make it easier to isolate when a generated trajectory has combined subskills from training trajectories.
VI-D1 Dataset Details
Table V describes details of the environment parameters and dataset description.
[…]
| Environment | Type | Embodiment | # of Trajectories | Trajectory Len | # of Obs | # of Acts | # of Possible Traj | # of Eval Traj |
| Maze | Numpy | Point | 316 | 1000 | 2 | 2 | 705,421 | 10000 |
| Didactic | Real | Franka | 3 | 12000 | 4 | 4 | 1 | 100 |
| Cube-Stack | RoboSuite | Franka | 100 | 1000 | 4 | 4 | 720 | 500 |
| Lights | OGBench | UR5e | 1000 | 500 | 4 | 4 | 65,536 | 2500 |
| Well-Plate Sim | RoboSuite | Franka | 50 | 1000 | 4 | 4 | 7056 | 500 |
| Well-Plate Real | Real | Franka | 50 | 51000 | 4 | 4 | 7056 | 100 |
VI-D2 Data Collection
Here we elaborate on our data collection process for each environment.
Maze: In Maze, we collect data by starting with the point agent in the top left corner. At each intersection, it randomly selects whether to travel down or to the right at that intersection, and repeats for future intersections. If the point agent is along the bottom or right-hand face, it will only go to the right or down respectively.
Well-Plate: In both the simulated and real Well-Plate environments, we enumerated all possibilities of start-plate cells to goal-plate cells (81, as there are 9 cells in each plate). Out of these 81 start-goal pairs, we randomly take 76 of them, and then randomly sample from the remaining pairs in groups of 3, excluding cases where there are collisions in either the start or goal plate. We did this until we had 50 trajectory descriptions consisting of three pick and places. To generate trajectories for these descriptions we generate a set of keypoints for the arm to travel between, and them follow them with a positional impedance controller. We fixed the RNG when generating the trajectory descriptions between the simulated and real well-plate environments, but the actual trajectories vary because of differences between the well plate placements between environments.
Block-Stack: In Block-Stack, to generate a trajectory description we index the coloured blocks from 0 to 5, generate all ordered permutations of those indexes, and then randomly select 100 from the list. To execute description we use a similar system to Well-Plate where we generate keypoints guiding the arm to pick up each block and place it at the stack location, advancing the height the place occurs at in accordance with the growing height of the stack.
Lights: We use the same trajectories in the puzzle-4x6-play-v0 dataset from Park et al. [64]. This environment has possible trajectories, as each button can either be turned on or off.
Didactic: For Didactic, we manually wrote the set of keypoints for the three training trajectories.
VI-D3 Environment Diversity Primitive Descriptions
Here we describe the environment observation primitives we use to describe diversity, as defined in Section III-E. We also describe how many possible unique trajectories exist under our primitive formulation. We do not define primitives for Didactic, because the task is designed to only have possible trajectory that completes the task.
Maze: In maze, the diversity primitives are the set of visited intersections in the maze. Intersections are the crossroads between the corners of four obstacles (or two obstacles and the wall). We calculate these visitations by taking the executed trajectory and seeing if the difference between each state in the trajectory and each intersection is smaller than some threshold. This means that each trajectory is defined by the set of intersections it passes through, and because each training trajectory moves from the top-left to bottom-right, each trajectory visits the same number of intersections. See Section VI-D5 for the number of possible trajectories.
Well-Plate: In well-plate, the diversity primitives are the set of occupied cells in the goal plate, and the empty cells in the starting plate. Each training trajectory moves three blocks from the starting plate to the goal place, so each training trajectory is defined by a set of primitives. This makes the diversity measure blind to the order the blocks are moved in, but the end result is still a large space of potential trajectories. The total number of possible complete trajectories is equal to , as you choose 3 blocks from the starting plate to move to 3 cells in the goal plate.
Block-Stack: In block-stack, the diversity primitives are tuples of a cube’s colour and it’s position along the stack. For instance, the set (Red,1),(Blue,2),(Yellow,3) is a possible trajectory description. The number of possible complete trajectories is , as you pick block from , then from the remaining , et cetera.
Lights: In Lights, the diversity primitives are binary variables associated with each light being turned on (via being pushed). The number of possible trajectories is , as each light can either be turned on or off.
VI-D4 Goal-Conditioned Environment Details
For the Goal-conditioned Maze experiment, we generated a new dataset in the Maze consisting of random walks from points along the edge of the maze to other edge points with trajectories. We then found the set of all edge point start-goal pairs unseen in the same training trajectory to use as evaluation pairs. This ensure that to complete the unseen goals, the diffusion model must stitch sub-trajectories from the training set without simply replicating existing trajectories.
VI-D5 Grid Environment Possible Trajectories
@@
VI-E Further Architecture Details and Experiments
VI-E1 U-Net is Not Shift Equivariant
@@
VI-E2 Engineering Details for Eq-Net
For an architecture diagram of Eq-Net, see Figure 15.
As noted in [42], in theory, circular padding is required for a CNN to have full shift equivariance, but in practice this is uncommon.
@@ VI-E2 Engineering Details for Eq-Net
VI-E3 Kernel Size Tuning
@@
Broadly, we found that tuning KER trades off between overall plan consistency (how likely the trajectories were to be locally consistent and stay outside the obstacles) and compositionality. To show this, we trained Eq-Net models in our unconditional setting with a variety of KER values. Our results are shown in Table VI. We found that the intermediate value of 10 generated both a strong diversity of plans as well as reliably well-composed plans, while making the model too non-local made the model too local and reduced completion rates.
[…]
| Task Completion % | 0.991 | 0.997 | 0.995 | 0.984 | 0.965 |
| % of Generations Novel | 0.5056 | 0.752 | 0.802 | 0.869 | 0.9701 |
Our experiments with the default U-Net architecture and our Eq-Net architectures in total confirm our earlier hypothesis that while CNNs are meant to have strong inductive biases to encourage composition, the particular set of decisions made by Janner et al. [40] means at small data amounts they can still overfit heavily and attend to non-local information.
VI-E4 Kernel Diversity
One thing that we found was important when designing Eq-Net was to include both very small (such as k=3) along with the larger kernel layers. When we used only medium sized kernels, it would lead to trajectories that were oddly "ungrounded" to the navigation plane, and would look correct in shape but would be shifted such that the resulting trajectory would nearly always hit obstacles. Figure 16 shows the visual results of this in early experiments in a maze. We found that increasing the kernel size from small to large during training helped, somewhat mirroring the process in normal CNNs where fixed-size kernels stride over the whole image initially (capturing fine detail) and then smaller subsets after each downstride (capturing coarse detail). However, we did not explore whether the exact ordering of the kernel layers in Eq-Net matters or not.
VI-E5 Positional Encoding
When testing how our small receptive field model performs when it has access to positional information for each state, we use the same positional encodings as in Vaswani et al. [75], with slightly varied parameters. Figure 17 shows a visualization of how we add positional encoding to the Eq-Net architecture. These encodings turn the position of each state into waves forming a geometric progression from which the position of a particular state can be inferred. The encodings are added after the first residual block. Specifically, the positional encodings are defined as
We used a maximum length of instead of used by Vaswani et al. [75] because of the shorter length of our sequences.
@@ VI-E5 Positional Encoding
VI-F More Experiments
VI-F1 Do Common Diffusion Model Parameters Impact Compositionality?
In the following sections we report the findings of exploratory experiments we performed on a wide host of other diffusion model parameters we examined. Ultimately, we found that none of the following factors have significant effects on composition, with the exception of overfitting which has a moderate effect. All these experiments were done in our Maze environment, with reported statistics taken over evaluation samples.
VI-F2 Does Using the Continuous or Discrete SDE Formulation Impact Compositionality?
No. In diffusion models, the discrete and continuous formulations refer to different ways to parameterize the underlying stochastic differential equation governing the diffusion process [60, 7]. Table VII shows that the continuous and discrete SDEs perform functionally identically.
[…]
| Completion % | 0.989 | 0.933 | 1 | 1 |
| % of Generations Novel | 0.815 | 0.7894 | 0 | 0 |
VI-F3 Does High-Temperature Sampling Impact Compositionality
Surprisingly, not substantially. We expected the diffusion sampling temperature to have some effect, but instead found that models behaved similarly regardless of temperature, with some enhanced compositionality only occurring at temperatures high enough to greatly degrade sample quality. Table VI-F3 shows that high-temperature sampling does not make any model more creative.
| Temperature | 0.1 | 0.5 | 1 | 2 | 4 |
[…]
| U-Net | 1 | 1 | 1 | 1 | 0 |
| DiT | 0.991 | 0.991 | 0.995 | 0.952 | 0 |
| Eq-Net | 0.989 | 0.987 | 0.989 | 0.993 | 0.26 |
[…]
| U-Net | 0 | 0 | 0 | 0 | NaN |
| DiT | 0.0226 | 0.0221 | 0.0189 | 0.0297 | NaN |
| Eq-Net | 0.827 | 0.801 | 0.806 | 0.815 | 0.796 |
VI-F4 Does Predicting Noise vs Clean Impact Compositionality?
No. In early experiments we found that predicting noise versus clean did not impact compositionality; however, it did impact sample quality (measured via completion rates). We found for the high-precision maze task that predicting noise worked better; when trained with the same parameters predicting a clean sample instead of noise, both U-Net and Eq-Net never successfully completed the maze. That our findings include environments where we train to predict noise (Maze) and where we train to predict clean (the manipulation ones) shows that the prediction target is likely irrelevant to compositionality.
VI-F5 Does the Number of Diffusion Sampling Steps Impact Compositionality?
Not past a certain, fairly low threshold. We found that while more sampling steps was very important to generate high-quality samples (measured by completion rate in our Maze environment, where slight perturbations from the maze causes the point agent to run into obstacles and die), we did not see a strong effect on compositionality, especially after roughly samples. Table VI-F5 shows the full results.
| # of Diffusion Sampling Steps | 10 | 50 | 100 | 500 | 1000 |
[…]
| U-Net | 0.839 | 1 | 1 | 1 | 1 |
| DiT | 0.409 | 0.914 | 0.973 | 0.994 | 0.995 |
| Eq-Net | 0.105 | 0.807 | 0.927 | 0.989 | 0.992 |
[…]
| U-Net | 0 | 0 | 0 | 0 | 0 |
| DiT | 0.009 | 0.00662 | 0.0258 | 0.009 | 0.032 |
| Eq-Net | 0.688 | 0.796 | 0.813 | 0.804 | 0.81 |
VI-F6 Does Overfitting Impact Compositionality?
@@
Yes. A common strategy in all machine learning to increase generalization is to perform early stopping, where training is ended before loss on the training set plateaus. Figure 18 shows that overfitting indeed reduces compositionality in the small-data setting, as models early in training achieve both high completion rates (indicating the generated plans are kinematically feasible) as well as more novel trajectories than later in training. However, the models that show stronger compositionality (such as Eq-Net) retain this property late into training (i.e, the losses in compositionality from overfitting plateaus). This hopefully indicates that the proper inductive biases from architecture choices can avoid the overfitting problem.
VI-G Expanded Main Text Experiment Results
VI-G1 Data Scaling Gradient Maps
Here we include the full gradient maps for the models described in Section IV-D. For spacing reasons, we include them at the very end of the appendix.
VI-H More Discussion
VI-H1 Broader Impacts